OpenStudy (anonymous):

find the critical points(solve for x) for the derivative of the function. f(x)= 5ln((x^2)+1)-3x

myininaya (myininaya):

\[f'(x)=5 \cdot \frac{(x^2+1)'}{(x^2+1)}-3\] you mean find the critical numbers of using f'

myininaya (myininaya):

you mean find the critical numbers of f using f'*

myininaya (myininaya):

\[5 \cdot \frac{2x}{x^2+1}-3=0\] solve for x

myininaya (myininaya):

\[\frac{10}{x^2+1}-\frac{3(x^2+1)}{x^2+1}=\frac{10-3x^2-3}{x^2+1}=\frac{-3x^2+7}{x^2+1}=0\] also f' exist every where so the only critical numbers are when \[-3x^2+7=0\]

myininaya (myininaya):

\[3x^2=7=> x^2=\frac{7}{3}=> x=\pm \sqrt{\frac{7}{3}}\]

OpenStudy (anonymous):

thanks you so much :)

myininaya (myininaya):

ninja i made a mistake

myininaya (myininaya):

\[\frac{10x}{x^2+1}-\frac{3(x^2+1)}{x^2+1}=\frac{10x-3x^2-3}{x^2+1}=\frac{-3x^2+10x-3}{x^2+1}=0\]

myininaya (myininaya):

\[-3x^2+10x-3=0 =>3x^2-10x+3=0=>3x^2-9x-x+3=0\] \[3x(x-3)-1(x-3)=(x-3)(3x-1)=0=> x=3 or x=\frac{1}{3}\]

OpenStudy (anonymous):

its cool, yeah that answer reflects the answers on the back of the book. thank you

myininaya (myininaya):

i left off my x when i went to the next step

myininaya (myininaya):

did you see that?

OpenStudy (anonymous):

yeah